Heat is added to 200 kg of ice at 0 ºC until dry steam at 100 ºC is obtained.

The specific heat capacity of water is 4200 J kg‑1 K‑1, the specific latent heat of steam is 2260 kJ kg-1 and the specific latent heat of ice is 335 kJ kg-1.

Determine the total amount of heat energy (in MJ) added to the ice.

a. 452MJ

b. 520MJ

c. 603MJ

d. 84MJ


1
Expert's answer
2021-07-06T11:37:05-0400

Gives

m=200kg

"L_{ic}=335kJkg^{-1}"

"L_{s}=2260KJkg^{-1}"

"s=4200Jkg^{-1}K^{-1}"

0° ice to 0°water

Q1=mL

"Q_1=200\times335=67000KJ"

0° water to100° water



"Q_2=ms∆T=200\times4200\times100=84000KJ"

100° water to100° steam

"Q_3=mL"

"Q_3=200\times2260=452000kJ"


"Q=Q_1+Q_2+Q_3=67000+84000+452000=603000KJ=603MJ"

option (c) is correct option